AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 14 of 33  ELOG logo
  ID Date Author Subject Text Attachments
  404   Fri May 26 08:53:54 2023 JM, CBStrain reliefs (flange side) installedStrain reliefs on the flange side of the
cable harness have been installed for two
detectors (top right and bottom right detectors). 
 strain-relief-flange-side.jpgstrain-relief1_(1).mp420230526_140923.jpgV-I-plot.png 
  403   Wed May 24 09:54:56 2023 JM, CBMaximum detector positionsThe detectors were moved into the beam
axis using the motors to check the maximum
distance they can be moved in without collision
 9x 
  402   Wed May 24 08:35:12 2023 JM, CBStrain reliefs (detector side) installedThe new MACOR strain reliefs on the (detector
side) have been installed. See attached photos

The two pieces  of MACOR are
 20230524_092010.jpg20230524_092017.jpg20230524_092101.jpg20230524_093000.jpg20230524_092159.mp4 
  401   Mon May 15 07:34:22 2023 TDMonday 15 May08.35 DSSSD#3 bias -149.95V leakage current
-6.418uA temperature +19.6 deg C ( https://web-docs.gsi.de/~lestinsk/tempplot.php
)
  
  400   Sun May 14 11:55:12 2023 TDSunday 14 May12.44 DSSSD#3 bias -149.95V leakage current
-6.448uA temperature +19.7 deg C ( https://web-docs.gsi.de/~lestinsk/tempplot.php
)
  
  399   Sat May 13 11:34:25 2023 TD, RSSSaturday 13 May
12.34 DSSSD#3 bias -149.95V leakage current
-6.763uA temperature +21.2 deg C ( https://web-docs.gsi.de/~lestinsk/tempplot.php
  
  398   Fri May 12 16:24:20 2023 RSS, TDV-I testSince yesterday evening, the bottom right
detector was at -150 V. In the morning, the
current was quite stable ~ 6.77 uA. The bias
 8x 
  397   Thu May 11 21:22:54 2023 RSS, TDV-I testToday, the foil was opened, the pins on
the bottom right detector were checked and
looked fine, the right-hand arm was moved
 7x 
  396   Wed May 10 19:06:18 2023 RSS, TDV-I testTo test the detectors, adaptors were installed
on the feed through flanges. A negative bias
voltage was applied in steps of 10 V from
 detector_1.pngdetector_2.pngdetector_3.pngdetector_4.png 
  395   Tue May 2 15:02:09 2023 CB, JMAll detectors and harnesses mounted Completed mounting harnesses and detectors.

The strain relief support for the
bottom most two detectors on the feedthrough
 20230502_153410.jpg20230502_153503.jpg20230502_153509.jpg20230502_153511.jpg 
  394   Mon May 1 18:21:18 2023 CB, JMAll detectors mounted Mounted four detectors on Carme ad well
as four cable harnesses.

 
 20230501_191533.jpg20230501_191511.jpg20230501_145740.jpg 
  393   Mon May 1 08:26:09 2023 CB, JMLeft hand detectors remounted (31 April)

Remounted 3335-12 on top and new
3335-1 on bottom of left hand plate. Used
  
  392   Sun Apr 30 10:01:39 2023 CB, JMLeft hand detectors unmounted (29 April)

Inspected CARME after opening.
See attached. Some key takeaways.
 26x 
  391   Sat Apr 29 10:12:58 2023 CB, JMCARME opened(28 April)

Mounted rails on dolly. The rails
don't smoothly connect to those on the
 8x 
  390   Thu Apr 27 19:43:23 2023 CB, JMPreparing to open Removed turbo pump and set aside in a box.
Considering switching to the other turbo
to avoid degradation due to lack of use.
 20230427_123456.jpg20230427_161602.jpg 
  389   Wed Apr 26 16:49:28 2023 CB, JMCARME dolly mostly built Disconnected cables from CARME motors and
turbo. The turbo has to go to allow Carme
to slide out. Crane may be required to lift
 20230426_152121.jpg20230426_162619.jpg20230426_162627.jpg 
  388   Tue Apr 25 13:22:19 2023 JMCARME venting CARME venting using nitrogen gas bottle. 

Coolant pipes dismounted from the
chamber
 20230425_134119.jpg20230425_132944.jpg 
  387   Mon Apr 24 21:24:26 2023 JMPreparation for CARME dismountCARME currently in vacuum, pressure ~3E-10
mbar. Chamber to be vented tomorrow afternoon
by vacuum division. We have active both pirani
 20230424_155635.jpg20230424_155607.jpg20230424_150526.jpg 
  386   Wed Mar 22 16:33:32 2023 TDNIM WR source
Test of NIM WR source at Edinburgh
 8x 
  385   Wed Jan 18 13:32:36 2023 PJCSMACB settings with either Emulator or VETARWhen using the VITAR connected
to the HDMI port of the root MACB the settings
for all the MACB in the system 0x3.
 macb_apr20.jedmacb_apr20.vhdzybo.jpgMACB.jpg 
ELOG V3.1.4-unknown